OpenStudy (anonymous):

please help Which ordered pair is the solution to the system of equations? x + y = -1 x - 3y = 11

OpenStudy (anonymous):

A. (2, −3) B. (−2, 1) C. (5, −2) D. (−6, 5)

OpenStudy (anonymous):

There is two ways you can get the answer. Solve the system, or put each choice into the equations to see if it works. I'm just going to solve it though, because it's much faster. :D x + y = -1 x - 3y = 11 I want to eliminate one variable to solve for the other. For this, I'm going to eliminate y by multiplying the top equation by 3. 3x + 3y = -3 x - 3y = 11 Now I can add them together. 4x + 0 = 8 4x = 8 Divide each side by 4. x = 2 <--- Now that we have this, we can plug it in to one of the original equations. x + y = -1 Plugging x in. 2 + y = -1 -2 -2 y = -3 So, the ordered pair that is a solution to the system of equations is: (2, -3) :D

OpenStudy (cggurumanjunath):

a

OpenStudy (anonymous):

ty :)

OpenStudy (anonymous):

You're welcome! :3